Let m represent the slope and b represent the y-intercept. Find the slope m and the y-intercept b for: 3x - 4y = 12.
Answer by usyim88hk(157)

3x - 4y = 12
We have to make y stand alone on one side to find m and b, so minus 3x from both sides
-4y = 12-3x
Divide both sides by -1
4y = 3x-12
divide both sides by 4
y = (3/4)x - 3
So m is (3/4) and b is -3